A test of hypotheses for one proportion has the following Null Hypothesis: pi = 0.33, and uses 0.10 for the level of significance. a. If the calculated value for the associated test statistic equals -1.71, determine the p-VALUE for the test Group of answer choices
Answer #1

To Test :-

H0 :- P = 0.33
H1 :- P ≠ 0.33

Test Statistic :-

Z = -1.71

P value = 2 * P ( Z > 1.71 ) = 2 * 1 - P ( Z < 1.71 ) = 0.0873 ( From Z table )


Reject null hypothesis if P value < α = 0.1 level of significance
Since 0.0873 < 0.1 ,hence we reject null hypothesis
Result :- Reject null hypothesis